# Work done during an internship at Meta FAIR

### v1

@@ Abstract
Figure 1: Short (average score across benchmarks) vs long context performance for 1.4B xLSTM, SWA (sliding window attention) and our selected SWAX design with different sliding window sizes.
@@ 1 Introduction
    We study the impact of the sliding window length on a wide range of tasks, encompassing validation perplexity, short-context reasoning, common sense benchmarks, and long-context modeling tasks;
@@
    We introduce a training strategy based on a stochastic window size. It provides, at the same time, the long-context performance enabled by short windows and the short-context and reasoning performance of longer windows.
@@ 2 Background
Equation (1) shows that if the kernel if $\phi$ is a finite-dimensional mapping, then the feature-mapped keys as well as the memory tensor are also finite-dimensional and can be materialized and cached for future retrievals $(\mathbf{H}_{t},\mathbf{z}_{t})$ in constant memory:
@@ 4.4 Different window sizes at train and test time
Figure 6: Average NIAH accuracy of 1.4B SWAX models depending on their train and test time window sizes.
@@
Figures 8 and 8 show that on RULER’s long-context recall tasks, the models trained with a stochastic window size and annealing perform on par or better than the model trained with the short window of 128, and drastically better than the hybrid model trained with a window of 2048 tokens. For instance, at 1.4B parameter scale depicted in Figure 8, the stochastic training SWAX model performs similarly to the short-window SWAX. Figure 8 shows that stochastic training also improves the long-context performance of models at 7B scale. At 7B scale, compared to using a fixed sliding window of size 2048, the stochastic training gives much better retrieval accuracy at all sequence lengths. Furthermore, just as at 1.4B scale, stochastic training gives similar or even superior long-context performance compared to using a short window throughout training. Overall, a stochastic window size at training time maintains all the benefits of having a short window for long-context recall, and most if not all of the benefits of a longer window for short/medium-context reasoning tasks.

### v2

@@ Abstract
Figure 1: Short (average score across benchmarks) vs long context performance for 1.4B xLSTM, SWA (sliding window attention) and SWAX with different sliding window sizes.
@@ 1 Introduction
    We investigate the impact of the sliding window length on a wide range of tasks, encompassing validation perplexity, short-context reasoning, common sense benchmarks, and long-context modeling tasks;
@@
    We present a training strategy based on a stochastic window size that provides the long-context performance enabled by short windows and the short-context and reasoning performance of longer windows.
@@ 2 Background
Equation (1) shows that if the kernel $\phi$ is a finite-dimensional mapping, then the feature-mapped keys as well as the memory tensor are also finite-dimensional and can be materialized and cached for future retrievals $(\mathbf{H}_{t},\mathbf{z}_{t})$ in constant memory:
@@ 4.4 Different window sizes at train and test time
Table 3: average NIAH accuracy of 1.4B SWAX models depending on their train and test time window sizes.
@@
Figures 7 and 7 show that on RULER’s long-context recall tasks, the models trained with a stochastic window size and annealing perform on par or better than the model trained with the short window of 128, and drastically better than the hybrid model trained with a window of 2048 tokens. For instance, at 1.4B parameter scale depicted in Figure 7, the stochastic training SWAX model performs similarly to the short-window SWAX. Figure 7 shows that stochastic training also improves the long-context performance of models at 7B scale. At 7B scale, compared to using a fixed sliding window of size 2048, the stochastic training gives much better retrieval accuracy at all sequence lengths. Furthermore, just as at 1.4B scale, stochastic training gives similar or even superior long-context performance compared to using a short window throughout training. Overall, a stochastic window size at training time maintains all the benefits of having a short window for long-context recall, and most if not all of the benefits of a longer window for short/medium-context reasoning tasks.
@@
More long-context benchmarks.
The Ruler tasks are artificial long context benchmarks, built to be sensible to model variants. We also evaluated SWAX on more realistic tasks from several families of benchmarks:
  • LongBench (Bai et al., 2024a) is a bilingual, multitask, and comprehensive benchmark to evaluate the long context understanding capabilities of LLMs. It is composed of six categories and twenty one different tasks. It evaluates long-text application scenarios such as single-document QA, multi-document QA, summarization, few-shot learning, synthetic tasks and code completion.
  • Babilong (Kuratov et al., 2024) This benchmark extends the Babi tasks (Weston et al., 2015), it is designed to evaluate long-term memory and reasoning capabilities of LLMs. Babilong uses long input sequences, making it suitable for evaluating models with advanced memory architectures. The main categories are: single book reasoning, memory and retrieval and temporal and spatial reasoning.
  • LongBench2 (Bai et al., 2024b) builds on the original LongBench, expanding both the scale and diversity of tasks to better stress-test LLMs for real-world long-context scenarios.
The results are summarized in Table 4 (bottom). They show that the tasks are difficult on average for such small models. Stochastic training outperforms fixed-size training in some settings, but for others, the long context (2048) performs better. In Table 6, Appendix D we present the average performance of the LongBench summarization and question-answering tasks, where the same conclusion holds.
Training SWA size
Task family 128 stochastic 2048
Longbench 10.67 11.79 11.14
Longbench2 22.99 27.26 22.65
Babilong 4.68 9.35 7.29
Gated DeltaNet
Ruler multiquery NIAH 40.58 40.90 28.80
Longbench 9.14 9.94 11.42
Babilong 3.74 7.21 8.95
Table 4: Other long-context benchmarks on our SWAX architecture with 1.4B parameters. The results are averaged over all tasks within a family. Top: xLSTM linear attention memory, bottom: Gated DeltaNet linear attention. For Gated DeltaNet we also report the averaged performance of Ruler multiquery NIAH, averaged over sequence sizes 4096, 8192 and 16384.
4.5 Experiments with Gated DeltaNet
The Gated DeltaNet (Yang et al., 2025) is a state-of-the-art hybrid architecture that combines gating and delta-update rule to update the linear attention memory. To assess whether our observations apply to other linear attention architectures, we replace the xLSTM component in the previous experiments with a Gated DeltaNet layer (we call this combination SWAX/GDN). We did not change any training setting, and ran the same long-context benchmarks. Table 4 (top) shows that hybrid architectures with Gated DeltaNets also benefit from the stochastic training.
@@ Appendix C Benchmarks
Appendix D LongBench summarisation and QA tasks performance
We can report the averages for all the summarization and question answering tasks, we observe that training with stochastic SWA size outperforms the fixed-size training windows.
Task SWAX:128 SWAX:2048 SWAX:stochastic
Longbench summarization tasks (average) 10.570 11.152 12.429
Longbench QA tasks (average) 5.203 6.112 7.257
Table 6: Average performance for summarization and question-answering LongBench tasks
  • The LongBench summarization tasks are: longbench_gov_report,longbench_qmsum, longbench_multi_news and longbench_vcsum.
  • The LongBench QA tasks, which include single-QA and multi-QA tasks, are: longbench_hotpotqa, longbench_2wikimqa, longbench_musique, longbench_dureader, longbench_narrativeqa, longbench_qasper, longbench_multifieldqa_en and longbench_multifieldqa_zh.